OpenStudy (anonymous):

how to find cos 6 theta?

OpenStudy (psymon):

Depends on what it equals to. cos(6x) = ?

OpenStudy (anonymous):

\[-1\le \Theta \le 1\] These are the only values accepted for sine and cosine, so like @Psymon said, it depends what its equal to.

OpenStudy (anonymous):

Okay.. I got it..

OpenStudy (psymon):

A lot of the time you want to solve simply for x first. For example, if we say cos(6x) = (1/2), we would first solve as if this were cosx = (1/2) So if wanted to find x, x would just be pi/3. So given that x = pi/3, in order to get 6x, we would write 6x = pi/3. In this case, dividing by 6 to get x by itself, x would be pi/18. This would be a normal way tosolve for x when you have something liek cos(6x)